# Chirality, confinement and dimensionality govern re-entrant transitions in active matter

### v1

@@ III Trapped and torque-driven ABP in 3D
It is straightforward to numerically integrate the above equations using the Euler-Maruyama scheme. As before, we can formulate the Fokker-Planck equation, apply the Laplace transform, and perform the subsequent analysis to obtain the various moments in steady state. Below, we summarize the key findings, with detailed derivations available in the supplementary material [71].
@@ III.1 Lower order moments at steady state
A symmetrized MSD in steady state reads
However, it is important to note that the torque choice results in rotational motion around the $z$-axis, confined to the xyx-y plane, while the dynamics along the $z$-axis exhibit a persistent random walk. To account for this dynamical asymmetry, we separately compute the component of MSD in the $xy$-plane,
@@
is independent of $\Omega$. Thus the MSD describes a prolate spheroid with $\langle{\tilde{\textbf{r}}}_{\perp}^{2}\rangle_{\rm st}<\langle{\tilde{\textbf{r}}}_{\parallel}^{2}\rangle_{\rm st}$ for any finite $\Omega$, $\langle{\tilde{\textbf{r}}}_{\perp}^{2}\rangle_{\rm st}$ shrinking with increasing $\Omega$.
III.2 Steady state excess kurtosis and phase diagram
@@
We use the exact analytic expression for the steady-state excess kurtosis $\tilde{\cal{K}}_{\rm st}$ (provided in the supplementary material [71]) to precisely quantify the non-Gaussian deviations. Figs. 3(a) and 3(b) display $\tilde{\cal{K}}_{\rm st}$ as a function of $\beta$ and $\Omega$, respectively. The solid lines represent the analytic results, which show excellent agreement with the simulation results depicted by the points.
@@
We plot $\tilde{\cal{K}}_{\rm st}$ as a function of $\beta$ for three $\Omega$ values $\Omega=0.1,~30,~10^{3}$ with ${\rm{Pe}}=10^{2}$ in Fig. 3(a). For an ABP under torque without a harmonic trap (or at small $\beta$ values), the long-time behavior shows zero excess kurtosis [56]. The steady state excess kurtosis initially deviates towards negative values (indicating activity-dominated behavior) as trap strength increases, before eventually returning to Gaussian behavior (zero excess kurtosis). This trend aligns with prior observations in two dimensions and with earlier studies of ABPs in a harmonic trap [66]. In contrast to two dimensions, in three dimensions the steady state excess kurtosis remains negative, indicating activity-dominated behavior in the high chirality regime. The behavior of the steady state excess kurtosis is qualitatively independent of chirality values because the persistence motion of particles is not restricted for high chirality, effectively displaying position distribution along the $z$-axis. We plot $\tilde{\cal{K}}_{\rm st}$ as a function of $\Omega$ for three $\beta$ values $\beta=0.1,~20,~2\times 10^{4}$ with ${\rm{Pe}}=10^{2}$ in Fig. 3(b). For small $\beta$ values, the behavior of the particle is nearly Gaussian (excess kurtosis $\tilde{\cal{K}}_{\rm st}\sim 0$) and independent of chirality. For $\beta=0.1$, the steady state excess kurtosis $\tilde{\cal{K}}_{\rm st}$ shows negative small values, indicating low activity regimes. For intermediate trap strength $\beta=20$, steady state excess kurtosis $\tilde{\cal{K}}_{\rm st}$ shows large negative values, indicating highly active regimes. The probability distribution of the position vector exhibits a bimodal form, with off-center peaks located at $\tilde{\rm{r}}^{ac}_{\parallel}={\rm{Pe}}/\beta$ along the $z$-axis and at $\tilde{\rm{r}}^{ac}_{\perp}={\rm{Pe}}/\sqrt{\beta^{2}+\Omega^{2}}$ in the xyx-y plane. For high value of trap strength $\beta=2\times 10^{4}$, steady state excess kurtosis $\tilde{\cal{K}}_{\rm st}$ shows the Gaussian behavior (nearly zero excess kurtosis), independent of chirality $\Omega$.
@@
Fig. 3(c) shows the steady state state diagram in the $\Omega-\beta$ plane, highlighting different states observed using excess kurtosis. The excess kurtosis $\tilde{\cal{K}}_{\rm st}$ quantify the active (negative excess kurtosis) and Gaussian passive (near-zero negligibly small negative excess kurtosis) region. For intermediate values of trap strength $\beta$, steady state excess kurtosis takes negative values around $\tilde{\cal{K}}_{\rm st}\sim-0.4$, known as the activity dominated regime. And, for small and large values of $\beta$, $\tilde{\cal{K}}_{\rm st}$ exhibits near zero values known as passive regime. The state diagram exhibiting re-entrant transition with $\beta$, similar to the behavior of ABP in a harmonic trap as presented in [66]. The excess kurtosis exhibits qualitatively similar behavior regardless of chirality values, unlike the two dimensional case where high chirality values leads to passive behavior discussed previously (see Fig. 2(c)). Thus, analyzing the position probability distribution in the $\tilde{\rm{r}}_{\perp}-\tilde{z}$ plane for varying chirality values $\Omega=0.1,~5,~30,~10^{2},~10^{3}$ at $\beta=20$, as shown in (d), is crucial for understanding the spatial behavior.
@@
Analysis of the Phase diagram: In contrast to the 2D case, the 3D scenario features two distinct active length scales, $\tilde{\rm{r}}_{\parallel}^{\rm ac}={\rm{Pe}}/\beta$, $\tilde{\rm{r}}_{\perp}^{\rm ac}={\rm{Pe}}/\sqrt{\beta^{2}+\Omega^{2}}$, defined along and perpendicular to the torque direction, respectively, adding complexity to the phase diagram. The active regime is bounded by three key conditions: (i) $\beta>1$, ensuring sufficient persistence relative to the trap relaxation time; (ii) $(\tilde{\rm{r}}_{\parallel}^{\rm ac})^{2}>1/\beta$, or equivalently $\beta<{\rm{Pe}}^{2}$, which guarantees that the off-center peak along the direction of torque exceeds thermal broadening; and (iii) $(\tilde{\rm{r}}_{\perp}^{\rm ac})^{2}>2/\beta$, corresponding to Eq.(8), ensuring appreciable off-center peak perpendicular to the torque. As shown in the analysis of Eq. (8), this final condition is well approximated by $\beta<{\rm{Pe}}^{2}/2$ and $\beta>2\Omega^{2}/{\rm{Pe}}^{2}$. The boundaries plotted in Fig.3(d) capture the key features of the phase diagram shown in Fig.3(c), which is derived from the full expression of excess kurtosis. The segment between points $(iii)$ and $(iv)$ in Fig. 3(c) marks a transition from a fully active regime to one where only the dynamics along the $z$-axis remain active, while motion in the $xy$-plane becomes effectively passive, characterized by trapped diffusion with an activity-dependent diffusivity. In 3D, $z$-dynamics remain active for $1<\beta<{\rm{Pe}}^{2}$ at all $\Omega$, precluding re-entrant transition.
@@
Figure 3(e) shows the evolution of the position distribution at fixed activity ${\rm{Pe}}=10^{2}$ and torque $\Omega=30$, as the trap strength $\beta$ is varied. The distribution exhibits approximate spherical symmetry in both the small- and large-$\beta$ limits. At low $\beta=0.04$, limited persistence results in equilibrium-like behavior with a Gaussian distribution. At the opposite extreme, for $\beta=2\times 10^{4}>{\rm{Pe}}^{2}$, thermal fluctuations dominate the dynamics in all directions, and the distribution again becomes Gaussian. In the intermediate regime, $5\leq\beta\leq 10^{3}$, activity plays a leading role, giving rise to distinctly non-Gaussian, anisotropic distributions. These are well described by the active length scales $\tilde{\rm{r}}_{\parallel}^{\rm ac}={\rm{Pe}}/\beta$ and $\tilde{\rm{r}}_{\perp}^{\rm ac}={\rm{Pe}}/\sqrt{\beta^{2}+\Omega^{2}}$. For smaller $\beta$ (e.g., $\beta=5<\Omega$), $\tilde{\rm{r}}_{\parallel}^{\rm ac}>\tilde{\rm{r}}_{\perp}^{\rm ac}$, producing a band-like distribution elongated along the $z$-axis. As $\beta$ increases, $\tilde{\rm{r}}_{\perp}^{\rm ac}$ grows and approaches $\tilde{\rm{r}}_{\parallel}^{\rm ac}$, gradually restoring isotropy. This crossover manifests in a transition from band-like to ring-like distributions.
@@ IV Conclusions
This framework reveals three characteristic regimes: negative kurtosis corresponding to bimodal, off-center active states; vanishing kurtosis indicating equilibrium-like passive behavior; and small positive kurtosis associated with weakly heavy-tailed distributions, a feature unique to two-dimensional chiral particles. The dimensionality of the system dictates whether re-entrant transitions survive or vanish. In two dimensions, increasing chirality suppresses the active phase and drives a crossover to passive behavior, while in three dimensions, torque preserves the active phase by maintaining directed motion along the torque axis, leading to anisotropic steady states. Simple active length-scale arguments capture the resulting phase boundaries and account for the observed shape of steady-state distributions.
Beyond their immediate relevance for trapped microswimmers, our predictions outline experimentally accessible signatures—such as kurtosis crossovers and anisotropic steady-state distributions—that can serve as benchmarks for artificial and biological realizations of chiral active matter. More broadly, our analysis demonstrates how confinement can be exploited to reveal hidden structure in active fluctuations, suggesting routes to engineer controllable steady states in chiral and torque-driven active matter.
@@ Acknowledgments
AP and AC thanks the computing facility at IISER Mohali. AC acknowledges support from the Indo-German grant (IC-12025(22)/1/2023-ICD-DBT). D.C. acknowledges C.Y. Cergy Paris Université for a Visiting Professorship, DAE (India) for grant 1603/2/2020/IoP/R&D-II/150288, and ICTS-TIFR, Bangalore, for an Associateship.

### v2

@@ III Trapped and torque-driven ABP in 3D
It is straightforward to numerically integrate the above equations using the Euler-Maruyama scheme. As before, we can formulate the Fokker-Planck equation, apply the Laplace transform, and perform the subsequent analysis to obtain the steady-state moments. Below, we summarize the key findings, with detailed derivations presented in the Appendix-C.
@@ III.1 Lower order moments at steady state
It is important to note that the chosen torque induces rotation around the $z$-axis, resulting in motion confined to the xyx-y plane, while dynamics along the $z$-axis follow a persistent random walk. This leads to dynamical asymmetry, which we address by separately computing the component of MSD in the $xy$-plane,
@@
with the last expression independent of $\Omega$. As expected, the dynamics along the $z$-axis agree with those of non-chiral ABPs in a harmonic trap [67]. Thus the MSD describes a prolate spheroid with $\langle{\tilde{\textbf{r}}}_{\perp}^{2}\rangle_{\rm st}<\langle{\tilde{\textbf{r}}}_{\parallel}^{2}\rangle_{\rm st}$ for any finite $\Omega$, $\langle{\tilde{\textbf{r}}}_{\perp}^{2}\rangle_{\rm st}$ shrinking with increasing $\Omega$.
A symmetrized MSD in the steady state, ${\langle}{\tilde{\textbf{r}}}^{2}{\rangle}_{\rm st}=\langle{\tilde{\textbf{r}}}_{\perp}^{2}\rangle_{\rm st}+\langle{\tilde{\textbf{r}}}_{\parallel}^{2}\rangle_{\rm st}$, reads
III.2 Steady-state excess kurtosis and phase diagram
@@
We use the exact analytic expression for the steady-state excess kurtosis $\tilde{\cal{K}}_{\rm st}$ (provided in the Appendix-C) to precisely quantify the non-Gaussian deviations. Figs. 3(a) and 3(b) display $\tilde{\cal{K}}_{\rm st}$ as a function of $\beta$ and $\Omega$, respectively. The solid lines represent the analytic results, which show excellent agreement with the simulation results depicted by the points.
@@
We plot $\tilde{\cal{K}}_{\rm st}$ as a function of $\beta$ for three $\Omega$ values $\Omega=0.1,~30,~10^{3}$ with ${\rm{Pe}}=10^{2}$ in Fig. 3(a). For an ABP under torque without a harmonic trap (or at small $\beta$ values), the long-time behavior shows zero excess kurtosis [56]. The steady-state excess kurtosis initially deviates towards negative values (indicating activity-dominated behavior) as trap strength increases, before eventually returning to Gaussian behavior (zero excess kurtosis). This trend aligns with prior observations in two dimensions and with earlier studies of ABPs in a harmonic trap [67]. In contrast to two dimensions, in three dimensions the steady state excess kurtosis remains negative, indicating activity-dominated behavior in the high chirality regime. The behavior of the steady state excess kurtosis is qualitatively independent of chirality values because the persistence motion of particles is not restricted for high chirality, effectively displaying position distribution along the $z$-axis. We plot $\tilde{\cal{K}}_{\rm st}$ as a function of $\Omega$ for three $\beta$ values $\beta=0.1,~20,~2\times 10^{4}$ with ${\rm{Pe}}=10^{2}$ in Fig. 3(b). For small $\beta$ values, the behavior of the particle is nearly Gaussian (excess kurtosis $\tilde{\cal{K}}_{\rm st}\sim 0$) and independent of chirality. For $\beta=0.1$, the steady-state excess kurtosis $\tilde{\cal{K}}_{\rm st}$ shows negative small values, indicating low activity regimes. For intermediate trap strength $\beta=20$, steady-state excess kurtosis $\tilde{\cal{K}}_{\rm st}$ shows large negative values, indicating highly active regimes. The probability distribution of the position vector exhibits a bimodal form, with off-center peaks located at $\tilde{\rm{r}}^{\rm ac}_{\parallel}={\rm{Pe}}/\beta$ along the $z$-axis and at $\tilde{\rm{r}}^{\rm ac}_{\perp}={\rm{Pe}}/\sqrt{\beta^{2}+\Omega^{2}}$ in the xyx-y plane. For high value of trap strength $\beta=2\times 10^{4}$, steady-state excess kurtosis $\tilde{\cal{K}}_{\rm st}$ shows the Gaussian behavior (nearly zero excess kurtosis), independent of chirality $\Omega$.
@@
Fig. 3(c) shows the steady-state phase diagram in the $\Omega-\beta$ plane, highlighting different states observed using excess kurtosis. The excess kurtosis $\tilde{\cal{K}}_{\rm st}$ quantify the bimodal active (negative excess kurtosis) and Gaussian-like (near-zero negligibly small negative excess kurtosis) region. For intermediate values of trap strength $\beta$, steady-state excess kurtosis takes negative values around $\tilde{\cal{K}}_{\rm st}\sim-0.4$, known as the activity dominated regime. And, for small and large values of $\beta$, $\tilde{\cal{K}}_{\rm st}$ exhibits near zero values of Gaussian-like regime. The phase diagram exhibiting re-entrant transition with $\beta$, similar to the behavior of ABP in a harmonic trap as presented in [67]. The excess kurtosis exhibits qualitatively similar behavior regardless of chirality values, unlike the two dimensional case where high chirality values leads to Gaussian-like behavior discussed previously (see Fig. 2(c)). Thus, analyzing the position probability distribution in the $\tilde{\rm{r}}_{\perp}-\tilde{z}$ plane for varying chirality values $\Omega=0.1,~5,~30,~10^{2},~10^{3}$ at $\beta=20$, as shown in (d), is crucial for understanding the spatial behavior.
@@
Analysis of the Phase diagram: In contrast to the 2D case, the 3D scenario features two distinct active length scales, $\tilde{\rm{r}}_{\parallel}^{\rm ac}={\rm{Pe}}/\beta$, $\tilde{\rm{r}}_{\perp}^{\rm ac}={\rm{Pe}}/\sqrt{\beta^{2}+\Omega^{2}}$, defined along and perpendicular to the torque direction, respectively, adding complexity to the phase diagram. The active regime is bounded by three key conditions: (i) $\beta>1$, ensuring sufficient persistence relative to the trap relaxation time; (ii) $(\tilde{\rm{r}}_{\parallel}^{\rm ac})^{2}>1/\beta$, or equivalently $\beta<{\rm{Pe}}^{2}$, which guarantees that the off-center peak along the direction of torque exceeds thermal broadening; and (iii) $(\tilde{\rm{r}}_{\perp}^{\rm ac})^{2}>2/\beta$, corresponding to Eq.(8), ensuring appreciable off-center peak perpendicular to the torque. As shown in the analysis of Eq. (8), this final condition is well approximated by $\beta<{\rm{Pe}}^{2}/2$ and $\beta>2\Omega^{2}/{\rm{Pe}}^{2}$. The boundaries plotted in Fig.3(d) capture the key features of the phase diagram shown in Fig.3(c), which is derived from the full expression of excess kurtosis. The segment between points $(iii)$ and $(iv)$ in Fig. 3(c) marks a transition from a fully active regime to one where only the dynamics along the $z$-axis remain active, while motion in the $xy$-plane becomes effectively Gaussian-like, characterized by trapped diffusion with an activity-dependent diffusivity. In 3D, $z$-dynamics remain active for $1<\beta<{\rm{Pe}}^{2}$ at all $\Omega$, thereby permitting a re-entrant transition with respect to $\beta$ across all values of $\Omega$.
@@
Figure 3(e) shows the evolution of the position distribution at fixed activity ${\rm{Pe}}=10^{2}$ and torque $\Omega=30$, as the trap strength $\beta$ is varied. The distribution exhibits approximate spherical symmetry in both the small- and large-$\beta$ limits. At low $\beta=0.04$, limited persistence results in equilibrium-like behavior with a Gaussian distribution. At the opposite extreme, for $\beta=2\times 10^{4}>{\rm{Pe}}^{2}$, thermal fluctuations dominate the dynamics in all directions, and the distribution again becomes Gaussian. In the intermediate regime, $5\leq\beta\leq 10^{3}$, activity plays a leading role, giving rise to distinctly non-Gaussian, anisotropic distributions. These are well described by the active length scales $\tilde{\rm{r}}_{\parallel}^{\rm ac}={\rm{Pe}}/\beta$ and $\tilde{\rm{r}}_{\perp}^{\rm ac}={\rm{Pe}}/\sqrt{\beta^{2}+\Omega^{2}}$. For smaller $\beta$ (e.g., $\beta=5<\Omega$), $\tilde{\rm{r}}_{\parallel}^{\rm ac}>\tilde{\rm{r}}_{\perp}^{\rm ac}$, producing a band-like distribution elongated along the $z$-axis. As $\beta$ increases, $\tilde{\rm{r}}_{\perp}^{\rm ac}$ approaches $\tilde{\rm{r}}_{\parallel}^{\rm ac}$, gradually restoring isotropy. This crossover manifests in a transition from band-like to ring-like distributions.
@@ IV Conclusions
This framework reveals three characteristic regimes: negative kurtosis corresponding to bimodal, off-center active states; vanishing kurtosis indicating Gaussian-like behavior; and small positive kurtosis associated with weakly heavy-tailed distributions, a feature unique to two-dimensional chiral particles. The dimensionality of the system dictates to what extent re-entrant transitions survive. In two dimensions, enhanced chirality diminishes non-Gaussian, bimodal features, promoting a transition toward Gaussian-like distributions. In three dimensions, torque preserves the non-Gaussian phase by sustaining trapped, non-chiral–like active motion along the torque axis, leading to anisotropic steady states. Simple active length-scale arguments capture the resulting phase boundaries and account for the observed shape of steady-state distributions.
Beyond their immediate relevance for trapped microswimmers, our predictions outline experimentally accessible signatures—such as kurtosis crossovers and anisotropic steady-state distributions — that can serve as benchmarks for artificial and biological realizations of chiral active matter. Our results are amenable to experimental verification in systems such as L-shaped colloidal swimmers [28], chiral granular rotors [78], and air-driven spinners [79]. More broadly, our analysis demonstrates how confinement can be exploited to reveal hidden structure in active fluctuations, suggesting routes to engineer controllable steady states in chiral and torque-driven active matter.
@@ Acknowledgments
AP and AC thanks the computing facility at IISER Mohali. AC acknowledges support from the Indo-German grant (IC-12025(22)/1/2023-ICD-DBT). D.C. acknowledges the Department of Atomic Energy (India) for grant 1603/2/2020/IoP/R&D-II/150288, CY Cergy Paris Université for a Visiting Professorship, and ICTS–TIFR, Bangalore, for an Associateship.